# PentestAgent: Incorporating LLM Agents to Automated Penetration Testing

### v1

@@ Abstract.
Recent advancements in large language models (LLMs) offer new opportunities for enhancing penetration testing through increased intelligence and automation. However, current LLM-based approaches still face significant challenges, including limited penetration testing knowledge and a lack of comprehensive automation capabilities. To address these gaps, we propose PentestAgent, a novel LLM-based automated penetration testing framework that leverages the power of LLMs and various LLM-based techniques like Retrieval Augmented Generation (RAG) to enhance penetration testing knowledge and automate various tasks. Our framework leverages multi-agent collaboration to automate intelligence gathering, vulnerability analysis, and exploitation stages, reducing manual intervention. We evaluate PentestAgent using a comprehensive benchmark, demonstrating superior performance in task completion and overall efficiency. This work significantly advances the practical applicability of automated penetration testing systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Penetration testingccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Although automated penetration testing has been a concept for over a decade, a significant gap remains between the proposed methods and their real-world application. Early works (Boddy et al., 2005; Obes et al., 2013; Roberts et al., 2011) primarily modeled attack planning as an attack graph problem (Ammann et al., 2002) in a deterministic and fully observable world. However, such an approach imposes limitations: it assumes complete observability from the defenders’ standpoint and lacks the flexibility and adaptability required for dynamic environments.
Later efforts (Durkota and Lisỳ, 2014; Krautsevich et al., 2013; Sarraute et al., 2011, 2013, 2012; Zhou et al., 2019; Hu et al., 2020; Chen et al., 2023) addressed these shortcomings by introducing uncertainty into planning methodologies, treating attack planning as a Markov Decision Process (MDP), which model the world as states and actions as transitions between states, with a reward function encoding the “reward” for moving from one state to another. As extensions to MDP-based approaches, the subsequent works employ partially observable Markov decision process (POMDP) (Sarraute et al., 2013, 2012) and reinforcement learning algorithms (Hu et al., 2020; Chen et al., 2023) to account for further uncertainty in the environment and action outcomes. These advancements better align with real-world conditions where attackers possess limited knowledge of the target systems. Nevertheless, these probabilistic models focus on establishing a theoretical model for automated pentesting planning and lack the implementation aspect.
@@
Without addressing these challenges, current methods become error-prone brute force techniques with limited effectiveness in practice.
@@
PentestAgent adopts a multi-agent design. Each agent within the framework is equipped with a set of tools and assumes responsibility for a specific task in the penetration testing process. These agents are highly adaptable, as their toolsets can be customized to suit various tasks, making them flexible and extensible across different scenarios.
Besides LLM agents, PentestAgent incorporates Retrieval Augmented Generation (RAG) (Lewis et al., 2020) into its framework. RAG enhances the capabilities of LLMs by enabling them to leverage additional data for response synthesis and thus serves as a potent augmentation tool for LLM agents, enabling them to produce more informed and contextually relevant outputs. RAG also allows customized control of the supplementary context in the communication, ensuring efficient use of the context window. Table  1 shows a comparison among existing automated pentesting systems.
PentestAgent comprises four major components: a reconnaissance agent, a search agent, a planning agent, and an execution agent. These components collaborate seamlessly to automate the three primary stages of penetration testing: intelligence gathering, vulnerability analysis, and exploitation.
The reconnaissance agent initiates the process by gathering environmental data upon receiving the target. It generates and executes reconnaissance commands to collect comprehensive information about the target host. This data is then analyzed and stored in an environmental information database for further reference.
In the vulnerability analysis stage, the search agent queries the environmental database to identify exposed services and applications. It identifies potential attack surfaces and procedures, cataloging them separately. Concurrently, the planning agent employs Retrieval Augmented Generation (RAG) techniques to refine potential attack surfaces and selects suitable exploits tailored to the target environment.
During exploitation, the execution agent attempts to execute planned attacks on the target host. It retrieves necessary operational details from the environmental database, debugs execution errors, and logs all activities for comprehensive penetration testing reports.
This comprehensive approach promises to mitigate the reliance on manual intervention and enhance the scalability and adaptability of automated pentesting systems. To sum up, we make the following contributions:
@@
We will make our benchmark datasets and framework publicly available to facilitate further research in automated penetration testing.
@@ 2.1. Penetration Testing
Penetration testing, or pentesting, is a multi-stage, labor-intensive process designed to identify security vulnerabilities in systems under test.
2.1.1. Workflow
According to the Penetration Testing Execution Standard (PTES) (Standard, 2024), penetration testing consists of three main stages: intelligence gathering, vulnerability analysis, and exploitation. Existing tools typically focus on individual tasks within these stages. For instance, Nmap(nmap, 2024) specializes in information gathering by collecting response data from a target through direct interaction. Nessus (Tenable, 2024) and OpenVAS (GreenBone, 2024) are dedicated to vulnerability analysis, providing comprehensive scanning capabilities through their integrated services and tools. Metasploit (Rapid7, 2024) focuses on exploitation, offering various exploits with customizable payloads once a vulnerability is identified. While these tools excel in their specific tasks, mastering their use and integrating them into a cohesive attack plan requires significant expertise in penetration testing and substantial manual effort.
@@ 2.1.1. Workflow
Recent advancements in artificial intelligence have led to the development of more sophisticated penetration testing frameworks based on machine learning and Markov Decision Process (MDP) algorithms (Zhou et al., 2019; Hu et al., 2020; Chen et al., 2023). For example, Chen et al. (Chen et al., 2023) designed a reinforcement learning-based framework for automated attack planning. This framework incorporates expert knowledge into state-action pairs and employs a reward function to train the system to execute actions with the highest success rate. Although these frameworks can generate reasonable attack plans, they lack the dynamic implementation aspects of penetration testing. They are unable to react to potential failures and adjust the plan in real time.
@@
The rise of LLM-based applications has further advanced the automation of penetration testing tasks such as text analysis, task planning, code modification, and execution debugging. However, the existing LLM-based penetration testing frameworks still lack comprehensive coverage of the stages and automation for practical use. AutoAttacker (Xu et al., 2024) focuses on constructing post-breach attacks, neglecting the pre-compromise stages. PentestGPT (Deng et al., 2024a), while implicitly considering multiple stages through its “pentesting task tree,” still relies on human decision to proceed with a certain branch of tasks, leading to inefficiency and ineffectiveness. For example, PentestGPT may overly focus on one task while neglecting others, resulting in an unbalanced approach. Moreover, PentestGPT and AutoAttacker depend on the LLM’s pre-trained knowledge and human analysis to gather additional information about the target, discover and validate vulnerabilities, and select the next steps from the task tree. These tasks still demand considerable manual effort.
@@
2.1.2. Scope
Depending on the scope, penetration testing can be roughly divided into two groups: external assessments and internal assessments. (Consulting, 2019) In external assessments, the penetration tester starts off on the internet and targets the client’s web applications, services, and other internet-facing assets. The tests include social engineering engagement, red team attack simulation, and external network compromise. In internal assessments, the tester has access to the internal network, source code, or physical access to a device under test. The tests include code review and internal network compromise. Several works have focused on improving internal assessments using LLM-based frameworks (e.g., ChatAFL (Meng et al., 2024) and FuzzGPT (Deng et al., 2024b) for fuzzing, LLift (Li et al., 2023b) and LATTE (Liu et al., 2023) for program and binary analysis), while the external assessments have gained less attention. This paper aims to fill the gap to improve external assessments with LLM-based methods. According to Rapid7’s latest report on penetration testing (Consulting, 2020), external assessments have been the most demanding ones, in which external network compromise alone takes more than 80% of the external assessments performed. Thus, this paper will illustrate the effectiveness of PentestAgent with web penetration testing.
@@ 2.2. Challenges of Applying LLM to Pentesting
Limited Pentesting Knowledge Example User: What may be the vulnerabilities of ActiveMQ 5.17.3? And how can I test if the vulnerabilities exist? GPT-4: (1) Keep Software Updated … (2) Check CVEs … (3) Security Scanning Tools … (4) Configuration Review … (5) Penetration Testing … (6) Log Analysis … User: Could you give detailed steps for pentesting? GPT-4: (1) Identify Known CVEs … (2) Set Up Metasploit … (3) Search for ActiveMQ Modules … (4) Configure and Run Exploit … (5) Manual Testing …
@@
Repetition of Tasks Example Intelligence Gathering LLM: Use Nmap to perform a comprehensive scan of all ports on the target host to identify open ports and services. User: {Nmap scan results} Vulnerability Analysis LLM: Use Nmap to perform a comprehensive scan of all ports on the target host to identify open ports and services. User: {Nmap scan results}
@@
Loss of Context Example Intelligence Gathering {Information collection steps}… LLM: The target OS is Linux and the target IP is 192.168.238.129. Exploitation User: How do I execute this exploit? LLM: The target OS and IP are needed to configure the exploit. For investigation of the unknown OS and IP, do the following: …
@@ 3.1. System Overview
Intelligence Gathering: 1 Upon receiving user input specifying the target, the reconnaissance agent initiates the penetration testing process by gathering environmental information about the target host. The reconnaissance agent generates and executes reconnaissance commands, aiming to collect comprehensive environmental data from the target host. 2 The reconnaissance agent then analyzes the execution results and compiles a summary of the target environment, which is stored in a designated environmental information database.
Vulnerability Analysis: Next, the search and planning agents work together to perform the vulnerability analysis. 3 The search agent queries the environmental information database to retrieve a list of services and applications exposed on the target host. 4 Guided by these services and applications, the search agent searches for potential attack surfaces and procedures and saves them in separate databases. 5 The planning agent first leverages the RAG techniques to find a list of potential attack surfaces. 6 Subsequently, the planning agent uses these identified attack surfaces to determine suitable exploits for the target environment.
Exploitation: 7 Finally, the execution agent attempts to execute these attack plans on the target host. 8 The execution agent communicates with the environmental information database to obtain the necessary information for executing the exploits. It also debugs any execution errors by modifying the code or executing additional commands to gather more information. 9 All execution history is stored in a database and can be used to generate a comprehensive penetration testing report.
@@ 3.2. Reconnaissance Agent
Reconnaissance System Message (Simplified) Role-play You’re an excellent cybersecurity penetration tester assistant. Guide the tester … Chain-of-Thought Use Nmap to identify exposed ports, then use relevant tools in Nmap to analyze these ports on the target host … RAG You should use your query tool to learn about available reconnaissance tools … Structured Output You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
After the reconnaissance agent determines that it should stop the reconnaissance loop, it summarizes the reconnaissance results and stores them in a database to make the short-term reconnaissance memory persistent. The following prompt generates a structured output of the reconnaissance summary. Specifying the output structure and providing a comprehensive example guides the agent to output relevant information and reduces hallucination.
@@ 3.3. Search Agent
Potential Attack Surface Analysis Prompt (Simplified) RAG & CoT Generate a concise summary of the document to answer the following questions: 1) Does this document describe vulnerabilities targeting a particular service or app; if so, what is the relevant service/app version? 2) Provide information that can be used to search for the exploit of the vulnerabilities. Structured Output You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} … For example, the response looks like this: {OUTPUT FORMAT EXAMPLE}
After analyzing all individual search results, the search agent summarizes them into a structured output for subsequent parsing and storing.
Search Results Summary Prompt List ALL CVE numbers, URLs, keywords, and their applicable version relevant to exploit the vulnerabilities of {APP}. The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
@@
Exploit Procedure Analysis Prompt (Simplified) RAG & CoT Give a concise summary of the entire repository to answer the following questions: 1) whether this repository contains an exploit targeting a particular service or app; 2) What effect does the exploit have? Use one phrase to summarize the effect (e.g., remote command execution); 3) What relevant service/app version can this exploit be applied to? Few-shot Learning Note the app version is typically formatted as x.y.z. Explicitly state the version with the following formats … 4) what are the requirements to run this exploit? (e.g., OS, library dependencies, etc.) Structured Output You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} … For example, the response looks like this: {OUTPUT FORMAT EXAMPLE}
@@
Figure 5. Hierarchical-tree structured pentesting knowledge database
@@ 3.4. Planning Agent
The planning agent takes detected services and applications from the reconnaissance agent as input and generates an exploitation plan as output. As shown in Fig. 1, the planning agent leverages RAG and the pentesting knowledge base (Fig. 5) to first generate a list of potential attack surfaces relevant to the services and applications. Then, the planning agent follows a similar process to generate a list of exploits.
The planning agent uses the service or application as a key to find the relevant database for potential attack surfaces and retrieves these from the database according to the service or application version and vulnerability types. The planning agent makes suggestions for attack surfaces based on the application version and categorizes the attack surfaces by vulnerability types. We designed the following prompt to generate a list of potential attack surfaces given a particular service or application.
Attack Surface Suggestion Prompt (Simplified) List out all vulnerabilities ranked by confidence that can be used to exploits … The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
The planning agent uses the attack surface to find the relevant database for exploits and retrieves exploit details from the database according to the service or application version and exploit effects (e.g., remote code execution, authentication bypass). The planning agent then makes suggestions for exploits based on the application version and categorizes the exploits by exploit effects. We designed the following prompt to generate a list of exploits for each potential attack surface.
Exploit Suggestion Prompt (Simplified) List out paths of all relevant repositories ranked by the confidence that contain exploits …The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
@@ 3.5. Execution Agent
The execution agent takes the exploit details as input and attempts to execute the exploits on the target automatically, ultimately generating an exploitation summary as output. The execution agent follows the order suggested by the planning agent. As illustrated in Fig. 6, each exploit execution can be divided into two stages: the preparation stage and the exploitation stage.
@@
During the preparation stage, the execution agent analyzes the exploit details to determine the requirements for successful execution, such as the parameters needed. It then queries the database containing environmental information to obtain the necessary information. The following system message guides the execution agent in fulfilling its duty. To bypass the safety mechanisms that may prevent the process while the execution agent tries to execute the exploit, we use role-playing to make it act as a cybersecurity penetration tester.
@@
Execution Preparation System Message (Simplified) Role-play You’re an excellent cybersecurity pentesting assistant … CoT Analyze the exploit to answer the following questions: 1) What parameters do you need to fill in to execute this exploit successfully? 2) What information do you need to fill in these parameters? RAG You should use your query tool to learn the details about the exploit … Structured Output You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
Execution Information Query Prompt (Simplified) Based on the known information, try to provide the information listed here. {INFO NEEDED …} CoT You should examine the information needed one by one. For each piece of information needed, you should … RAG You should use your query tool to learn about the target environment … Structured Output The results should be presented in valid JSON format with the following fields: {FORMAT SPEC.} …
After getting the response containing the requested information, the execution agent’s system message is updated to transition to the exploitation stage.
Execution Exploit System Message (Simplified) Your next task is to provide step by step guide for executing the exploit and debugging the errors encountered … RAG You should use the tool to learn the code and README of the exploit to figure out how to properly execute it. Specifications Avoid commands that require user interactions … Self-reflection When the results indicate an error, you should … Structured Output You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …

### v3

@@ Abstract.
Recent advancements in large language models (LLMs) offer new opportunities for enhancing penetration testing through increased intelligence and automation. However, current LLM-based approaches still face significant challenges, including limited penetration testing knowledge and a lack of comprehensive automation capabilities. To address these gaps, we propose PentestAgent, a novel LLM-based automated penetration testing framework that leverages the power of LLMs and various LLM-based techniques like retrieval augmented generation (RAG) to enhance penetration testing knowledge and automate various tasks. Our framework leverages multi-agent collaboration to automate intelligence gathering, vulnerability analysis, and exploitation stages, reducing manual intervention. We evaluate PentestAgent using a comprehensive benchmark, demonstrating superior performance in task completion and overall efficiency.
@@
copyright: acmlicensedjournalyear: 2025copyright: ccconference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Hanoi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Hanoi, Vietnamdoi: 10.1145/3708821.3733882isbn: 979-8-4007-1410-8/2025/08ccs: Security and privacy Penetration testingccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Although automated penetration testing has been a concept for over a decade, a significant gap remains between the proposed methods and their real-world application. Early works (Boddy et al., 2005; Obes et al., 2013; Roberts et al., 2011) primarily modeled attack planning as an attack graph problem (Ammann et al., 2002) in a deterministic and fully observable world. However, such an approach imposes limitations: it assumes complete observability from the defenders’ standpoint and lacks the flexibility and adaptability required for dynamic environments. Later efforts (Durkota and Lisỳ, 2014; Krautsevich et al., 2013; Sarraute et al., 2011, 2013, 2012; Zhou et al., 2019; Hu et al., 2020; Chen et al., 2023) addressed these shortcomings by introducing uncertainty into planning methodologies, treating attack planning as a Markov Decision Process (MDP), which model the world as states and actions as transitions between states, with a reward function encoding the “reward” for moving from one state to another. As extensions to MDP-based approaches, the subsequent works employ partially observable Markov decision process (POMDP) (Sarraute et al., 2013, 2012) and reinforcement learning algorithms (Hu et al., 2020; Chen et al., 2023) to account for further uncertainty in the environment and action outcomes. These advancements better align with real-world conditions where attackers possess limited knowledge of the target systems. Nevertheless, these probabilistic models focus on establishing a theoretical model for automated pentesting planning and lack the implementation aspect.
@@
PentestAgent employs a multi-agent design where each agent is responsible for a specific task in the penetration testing process. This flexible architecture enables the customization of toolsets across different tasks, enhancing the system’s adaptability. In addition to LLM agents, PentestAgent integrates Retrieval Augmented Generation (RAG) (Lewis et al., 2020) to leverage supplementary data during response synthesis and manage the context efficiently. This combination enhances the quality of the generated outputs and further reduces the need for manual intervention.
Table 1 provides a comparative overview of existing automated pentesting systems, illustrating that while frameworks like AutoAttacker and PentestGPT rely on manual or limited approaches, PentestAgent offers a fully automated solution with a larger state and action space and advanced online search augmentation and debugging capabilities.
Moreover, our contributions extend beyond the framework itself. We introduce a comprehensive penetration testing benchmark based on VulHub’s collection of vulnerable Docker environments and Capture The Flag (CTF) challenges on HackTheBox. This benchmark spans various difficulty levels and encompasses a wide range of vulnerabilities, addressing a critical gap in current research by providing a practical and accessible evaluation framework.
To sum up, we make the following contributions:
@@
We make our benchmark datasets and framework publicly available to facilitate further research in automated penetration testing. [^1] https://github.com/nbshenxm/pentest-agent
@@ 2.1. Penetration Testing
Penetration testing (pentesting) is a structured, multi-stage process designed to identify security vulnerabilities in systems. According to the Penetration Testing Execution Standard (PTES) (Standard, 2024), pentesting consists of three main stages: intelligence gathering, vulnerability analysis, and exploitation. Penetration testing is broadly divided into external and internal assessments (Consulting, 2019). External assessments focus on assets exposed on internet, such as web applications, online services, and external networks, using techniques like social engineering, red teaming, and, in particular, web penetration testing. In contrast, internal assessments target the organization’s internal network, source code, or physical devices, typically involving code reviews and internal network compromises.
While several recent works have enhanced internal assessments using LLM-based frameworks (e.g., ChatAFL (Meng et al., 2024), FuzzGPT (Deng et al., 2024b), LLift (Li et al., 2023b), and LATTE (Liu et al., 2023)), external assessments, especially web penetration testing, remain underexplored. According to Rapid7’s latest report (Consulting, 2020), external network compromise constitutes over 80% of penetration testing tasks. This paper addresses this gap by demonstrating how PentestAgent automates web penetration testing, thereby enhancing the overall applicability and efficiency of automated pentesting in real-world scenarios.
Existing Tools. While automated tools exist for individual tasks within these stages, integrating them into a seamless and effective workflow remains a significant challenge. Existing tools specialize in specific penetration testing tasks. For instance, Nmap(nmap, 2024) is widely used for intelligence gathering, allowing testers to analyze network configurations through direct interaction with targets. Nessus(Tenable, 2024) and OpenVAS(GreenBone, 2024) focus on vulnerability analysis, scanning systems for known weaknesses using extensive vulnerability databases. Metasploit(Rapid7, 2024) is commonly used for exploitation, providing a range of exploits and payloads to execute attacks on identified vulnerabilities. Although these tools are effective alone, their effective use requires expert knowledge, manual decision making, and significant effort to coordinate workflow.
@@
AI-Driven Penetration Testing. Recent advancements in artificial intelligence have led to the development of more sophisticated penetration testing frameworks based on machine learning and Markov Decision Process (MDP) algorithms (Zhou et al., 2019; Hu et al., 2020; Chen et al., 2023). For example, Chen et al. (Chen et al., 2023) designed a reinforcement learning-based framework for automated attack planning. This framework incorporates expert knowledge into state-action pairs and employs a reward function to train the system to execute actions with the highest success rate. Although these frameworks can generate reasonable attack plans, they lack the dynamic implementation aspects of penetration testing. They are unable to react to potential failures and adjust the plan in real time.
@@
LLM-Based Penetration Testing. The rise of LLM-based applications has further advanced the automation of penetration testing tasks such as text analysis, task planning, code modification, and execution debugging. However, the existing LLM-based penetration testing frameworks still lack comprehensive coverage of the stages and automation for practical use. AutoAttacker (Xu et al., 2024) focuses on constructing post-breach attacks, neglecting the pre-compromise stages. PentestGPT (Deng et al., 2024a), while implicitly considering multiple stages through its “pentesting task tree,” still relies on human decision to proceed with a certain branch of tasks, leading to inefficiency and ineffectiveness. For example, PentestGPT may overly focus on one task while neglecting others, resulting in an unbalanced approach. Moreover, PentestGPT and AutoAttacker depend on the LLM’s pre-trained knowledge and human analysis to gather additional information about the target, discover and validate vulnerabilities, and select the next steps from the task tree. These tasks still demand considerable manual effort.
@@ 2.2. Challenges of Applying LLM to Pentesting
Limited Pentesting Knowledge Example
User: What may be the vulnerabilities of ActiveMQ 5.17.3? And how can I test if the vulnerabilities exist?
GPT-4:
  1. (1)
    Keep Software Updated …
  2. (2)
    Check CVEs …
  3. (3)
    Security Scanning Tools …
  4. (4)
    Configuration Review …
  5. (5)
    Penetration Testing …
  6. (6)
    Log Analysis …
User: Could you give detailed steps for pentesting?
GPT-4:
  1. (1)
    Identify Known CVEs …
  2. (2)
    Set Up Metasploit …
  3. (3)
    Search for ActiveMQ Modules …
  4. (4)
    Configure and Run Exploit …
  5. (5)
    Manual Testing …
@@
Repetition of Tasks Example
Intelligence Gathering
LLM: Use Nmap to perform a comprehensive scan of all ports on the target host to identify open ports and services.
User: {Nmap scan results}
Vulnerability Analysis
LLM: Use Nmap to perform a comprehensive scan of all ports on the target host to identify open ports and services.
@@
Loss of Context Example
Intelligence Gathering
{Information collection steps}…
LLM: The target OS is Linux and the target IP is 192.168.238.129.
Exploitation
User: How do I execute this exploit?
LLM: The target OS and IP are needed to configure the exploit. For investigation of the unknown OS and IP, do the following: …
@@ 3.1. System Overview
[…]
@@ 3.2. Reconnaissance Agent
Reconnaissance System Message (Simplified)
Role-play
You’re an excellent cybersecurity penetration tester assistant. Guide the tester …
Chain-of-Thought
Use Nmap to identify exposed ports, then use relevant tools in Nmap to analyze these ports on the target host …
RAG
You should use your query tool to learn about available reconnaissance tools …
Structured Output
You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
@@
After the reconnaissance agent determines that it should stop the reconnaissance loop, it summarizes the reconnaissance results and stores them in a database to make the short-term reconnaissance memory persistent.
@@ 3.3. Search Agent
Potential Attack Surface Analysis Prompt (Simplified)
RAG & CoT
Generate a concise summary of the document to answer the following questions:
1) Does this document describe vulnerabilities targeting a particular service or app; if so, what is the relevant service/app version?
2) Provide information that can be used to search for the exploit of the vulnerabilities …
Structured Output
You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
For example, the response looks like this: {OUTPUT FORMAT EXAMPLE}
@@
Figure 5. Hierarchical pentesting knowledge database
@@ 3.4. Planning Agent
The planning agent takes the detected services and applications from the reconnaissance agent as input and generates an exploitation plan as output. As shown in Fig. 1, the planning agent leverages RAG and the pentesting knowledge base (Fig. 5) to first generate a list of potential attack surfaces relevant to services and applications. Then, the planning agent follows a similar process to generate a list of exploits.
The planning agent uses the service or application as a key to find the relevant database for potential attack surfaces and retrieves these from the database according to the version of the service or application and the types of vulnerabilities. The planning agent makes suggestions for attack surfaces based on the application version and categorizes attack surfaces by vulnerability types.
The planning agent then uses the attack surface to find the relevant database for exploits and retrieves exploit details from the database according to the service or application version and exploit effects (e.g., remote code execution, authentication bypass). The planning agent then makes suggestions for exploits based on the application version and categorizes the exploits by exploit effects.
@@ 3.5. Execution Agent
The execution agent takes the details of the exploit as input and attempts to execute the exploits on the target automatically, ultimately generating an exploitation summary as output. The execution agent follows the order suggested by the planning agent. As illustrated in Fig. 6, each exploit execution can be divided into two stages: the preparation stage and the exploitation stage.
@@
During the preparation stage, the execution agent analyzes the details of the exploit to determine the requirements for successful execution, such as the parameters needed. It then queries the database containing environmental information to obtain the necessary information. The following system message guides the execution agent in fulfilling its duty. To bypass the safety mechanisms that may prevent the process while the execution agent tries to execute the exploit, we use role-playing to make it act as a cybersecurity penetration tester.
@@
Execution Preparation System Message (Simplified)
Role-play
You’re an excellent cybersecurity pentesting assistant …
CoT
Analyze the exploit to answer the following questions:
1) What parameters for executing this exploit?
2) What information do you need to execute the exploit?
RAG
Use your query tool to learn the details about …
Structured Output
You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …
After obtaining a list of needed information, the execution agent queries the environmental information database to fill in the information. After receiving the response containing the requested information, the execution agent’s system message is updated as follows to transition to the exploitation stage.
Execution Exploit System Message (Simplified)
  Your next task is to provide step by step guide for executing the exploit and debugging the errors encountered …
RAG
You should use the query tool to learn the code and README of the exploit to figure out how to properly execute it. You also use …
Self-reflection
When the results indicate an error, you should analyze the error and try to fix it …
Structured Output
You should always respond in valid JSON format with the following fields: {FORMAT SPEC.} …